OpenStudy (anonymous):

Which function is vertically stretched by a factor of 6 and translated 8 units right from the parent function?

OpenStudy (anonymous):

@SolomonZelman Could you help me with this problem as well please?

OpenStudy (solomonzelman):

general translations from f(x)=2x f(x)=2(x+1) 1 unit left f(x)=2(x-1) 1 unit right f(x)=2x +1 1 unit up f(x)=2x -1 1 unit down and f(x)=6x is going to make it 3 times steeper.

OpenStudy (anonymous):

okay, where does the 8 units right play a role in this? @SolomonZelman

OpenStudy (solomonzelman):

What is your parent function ?

OpenStudy (anonymous):

it didn't give us one

OpenStudy (solomonzelman):

I would assume that the parent is f(x)=x a plain y=x, then it would be f(x)=6(x-8)

OpenStudy (solomonzelman):

I think.....

OpenStudy (anonymous):

the correct answer ia 6log(x-8)

OpenStudy (solomonzelman):

Well.... "log" I am sorry... but how was I supposed to know that? I think that you were given that, or otherwise it is an unfair question.

OpenStudy (anonymous):

that's what the teacher gave us - i guess it's her fault then i don't know

OpenStudy (solomonzelman):

Well, she have probably given you the parent function...idk -:(